# Massive scalar perturbations in Kerr Black Holes: near extremal analysis

### v2

@@ II Scalar perturbations in Kerr Black Holes
$\partial_{u}\left[(1-u^{2})\partial_{u}S\right]+\left(\frac{m^{2}}{1-u^{2}}-a^{2}c^{2}u^{2}-\Lambda\right)S=0,$ (6)
where $u=\cos\theta$, $c=a\sqrt{\omega^{2}-\mu^{2}}$, and $\Lambda=\lambda+2am\omega-a^{2}\omega^{2}$.
@@
$R(r)\rightarrow\begin{cases}e^{-i(\omega-m\Omega_{+})r_{*}},\ \ \ r\rightarrow r_{+}\ \ (r_{*}\rightarrow-\infty),\\ \frac{1}{r}e^{+\sqrt{\mu^{2}-\omega^{2}}r_{*}},\ \ \ r\rightarrow\infty\ \ (r_{*}\rightarrow\infty),\end{cases}$ (7)
@@ III.1 Boundary conditions and monodromy parameters
We note that $y_{t_{0},\pm}$ and $y_{\infty,\pm}$ form a basis of solutions near $z=t_{0}$ and $z=\infty$, respectively.
@@
$\mathsf{C}_{t}=\begin{pmatrix}e^{-\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{t_{0}}-e^{\tfrac{i\pi}{2}\eta}\zeta_{t_{0}}&-e^{-\tfrac{i\pi}{2}\eta}\zeta_{\infty}\zeta^{\prime}_{t_{0}}+e^{\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{\infty}\zeta_{t_{0}}\\ e^{-\tfrac{i\pi}{2}\eta}-e^{\tfrac{i\pi}{2}\eta}&-e^{-\tfrac{i\pi}{2}\eta}\zeta_{\infty}+e^{\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{\infty}\end{pmatrix}$ (14)
@@
$\zeta_{\infty}=e^{\frac{i\pi}{2}\sigma}\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma),$ (15a)
$\zeta^{\prime}_{\infty}=e^{-\frac{i\pi}{2}\sigma}\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma),$ (15b)
@@
The QNM problem requires solutions of the radial equation to be purely ingoing at $z=t_{0}$ and purely outgoing at $z=\infty$. This translates into $\mathsf{C}_{t}$ being a lower triangular matrix, and implies the following relation between the monodromy parameters
$e^{i\pi\eta}=\frac{\zeta_{\infty}\zeta^{\prime}_{t_{0}}}{\zeta^{\prime}_{\infty}\zeta_{t_{0}}}=e^{i\pi\sigma}\frac{\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma)}\frac{\sin\tfrac{\pi}{2}(\theta_{t_{0}}+\theta_{0}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{t_{0}}+\theta_{0}-\sigma)}$
@@
We remark that $\rho$ and $\nu$ are canonically conjugate variables, in the sense that $\rho=(\pi i)^{-1}\partial_{\nu}{\cal W}$, just like the relationship between $\eta$ as $\sigma$ given in Eq. (22). Knowledge of $\eta$, together with $\sigma$, is required to obtain the transformation from $\{\sigma,\eta\}$ to $\{\nu,\rho\}$, allowing us to compute the accessory parameter $c_{t_{0}}$ from two different expansions, one in powers of $t_{0}$ and another in powers of $t_{0}^{-1}$. This provides not only for an independent numerical check of the eigenvalue calculation but also better behavior of the equations near extremality.
@@ III.2 Accessory parameter expansions via continued fractions
To compute the accessory parameter $c_{t_{0}}$ we employ a technique based on continued fractions. We consider Floquet type solutions of the CHE (9[72],
@@ III.3 The isomonodromic method for the QNM problem
$\theta_{0}=-m,\quad\theta_{t_{0}}=m,\quad\theta_{*}=0,\quad t_{0}=-2\alpha$ (38a)
$t_{0}c_{t_{0}}=\lambda+(1-m)\alpha+\frac{\alpha^{2}}{4},$ (38b)
where
$\alpha=2a\sqrt{\omega^{2}-\mu^{2}},$ (39)
@@
For small $\alpha$, we can derive an analytic expression for $\lambda=\lambda(\omega)$ by solving (41). Indeed, using the expansion (31) in powers of $t_{0}$, we find
$\begin{gathered}\lambda(\alpha)=\ell(\ell+1)+\frac{1}{4}\bigg{(}\frac{2((\ell+1)^{2}-m^{2})(\ell+1)^{4}}{(2\ell+1)(\ell+1)^{3}(2\ell+3)}\\ \qquad\qquad-\frac{2(\ell^{2}-m^{2})\ell^{4}}{(2\ell-1)\ell^{3}(2\ell+1)}-1\bigg{)}\alpha^{2}+{\mathcal{O}}(\alpha^{3}).\end{gathered}$ (42)
For large $\alpha$, a similar expression can be constructed from (36), with a suitable choice of $\nu$. In fact, the expression can be easily obtained from Eq. (4.7) in [41] by simply changing $a\omega$ to $\alpha/2$. Since the expansion is not necessary for the remainder of this paper, we will leave out its explicit form.
@@
$z=2i\omega(r-r_{-}).$ (43b)
The single monodromy parameters $\{\theta_{k}\}$ are expressed in terms of the physical parameters through
@@
$\theta_{\star}=\frac{2M(\mu^{2}-2\omega^{2})}{\sqrt{\mu^{2}-\omega^{2}}}.$ (44c)
@@
$\sqrt{\mu^{2}-\omega^{2}}(r_{-}-r_{+}-2iam+2i(a^{2}+r_{+}^{2})\omega)$
$+\frac{M(\mu^{2}-2\omega^{2})}{\sqrt{\mu^{2}-\omega^{2}}}-\frac{M(\mu^{2}-2\omega^{2})^{2}}{2(\mu^{2}-\omega^{2})}.$
@@
$g_{1}(\sigma,\eta,\omega)=0.$ (46)
Additionally, substituting Eqs. (44) and (45) into the continued fraction equation (30), yields an implicit relation, denoted by $g_{2}$, between $\sigma$, $\omega$, and $\lambda$,
@@
Hence, in order determine the QNM frequencies through the isomonodromic method we need to determine the numerical solutions of the coupled system of algebraic equations (41), (46)-(48). In practice, we fix the black hole parameters $M$ and $a$ and the scalar field parameters $\mu$, $\ell$, $m$ and then determine the QNM frequency $\omega$ (and also the associated parameters $\lambda$, $\sigma$, and $\eta$). As explained before, there exists an infinite set of solutions $\omega_{n}$ indexed by the non-negative integer $n$. Our numerical routine uses the implementation of $\tau_{V}$ as a Fredholm determinant truncated at $N_{f}=64$ Fourier components. The continued fractions are truncated at $N_{c}=128$ and the root-finding algorithm we use is Muller’s method. The interested reader can find the details of our implementation of the isomonodromic method, based on the Julia language, in [38, 77].
@@ IV.1 QNMs for $0<m<\ell$
Note that both the imaginary and the real parts of the QNM fundamental frequency converge to a finite value. This finite value is the QNM frequency of the extremal black hole, computed using the RH map for the $\tau_{III}$-function defined in [38]. The $\tau_{III}$-function must be used instead of the $\tau_{V}$-function because the radial equation is doubly confluent when the black hole is extremal. We thus have found strong numerical evidence that the massive scalar QNM frequencies are continuous functions of the spin $a/M$ when $a/M=1$, just like the corresponding massless frequencies [48, 38].
@@ V Zero-damping modes (ZDMs): Analytic expansion near extremality
in the near extremal regime. This expression mirrors the ones deduced in the literature for massive [82] and massless [83, 84, 38] scalar fields around near extremal black holes – see also Refs. [23, 85, 86, 87, 47, 88].
@@
We close this section with a discussion about the choice for the sign of $\sigma_{0}$ in Eq. (54a), which also determines the choice of the sign in the analytic expression (59) for the ZDMs near extremality. If $\sigma_{0}$ is real, which is expected for small $m$, we recall the analysis below Eq. (56) to justify that the signal of $\sigma_{0}$ must chosen to enforce $\mathrm{Re}(\sigma_{0})>0$, which selects the positive root in (54) This seems to be the case only for the $\ell=m=1$ ZDM, as illustrated in Fig. 6, where we observe that the correction is of order $\mathcal{O}(\delta^{2},\delta\log\delta)$ to the real part of the frequencies $\omega_{k}$ for the first three ZDMs ($k=0$, $k=1$, and $k=2$). In particular, we see in the plots that the values of $\mathrm{Re}(M\omega_{k})$ for $k=1$ and $k=2$ intercept the corresponding value for the $k=0$ mode at $\delta\simeq 2.5\times 10^{-2}$ and at $\delta\simeq 1.5\times 10^{-2}$ respectively. For the imaginary part of the frequency $\mathrm{Im}(M\omega_{k})$, the linear $\delta$ term in (59) describes well the near extremal behaviour presented in Fig. 6. In Table 1, we compare the QNMs obtained through the numerical method detailed in Sec. III.3 with the frequencies calculated using the asymptotic expansion (59).
@@
Figure 7: Near extremal behavior of the first three ZDMs when $\ell=m=2$. The left and right panels show, respectively, the real and imaginary parts of the QNM frequencies for selected values of the scalar field mass.
@@ VI Continued fraction for large $t_{0}$ and contour plots
which is the continued fraction equation after the elimination of the dependence on the variables $\nu$ and $\lambda$. In order gain insights on the QNM spectrum, we generate contour plots of the function $G(\omega)$ for complex values of the argument.
We start by checking whether the QNM frequencies calculated by finding the roots of $G$, which is based on large $t_{0}$ expansions, agrees with the results shown in Fig. 1 for $\ell=1$ and $m=0$, which were obtained with the isomonodromic method of Sec. III.3 that is based on power series of $t_{0}$. Table 2 presents the fundamental QNM frequency obtained by solving Eq. (60) for several masses $M\mu$ and spins $a/M$ when $\ell=1$ and $m=0$. We have confirmed that the values obtained from Fig. 1 are consistent with those listed in Table 2, with an agreement to order $10^{-10}$. Being DMs, as $a/M\rightarrow 1$, the real and imaginary parts of their frequencies converge to a finite value.
@@
We had previously identified the emergence of a bifurcation point for the $\ell=m=1$ fundamental QNM in the near extremal regime, as shown in Figs. 4 and 5. Beyond this bifurcation point, i.e. for $M\mu>(M\mu)_{c}\simeq 0.37049$ and $a/M>(a/M)_{c}\simeq 0.99946$, the $n=0$ QNM is a DM and not a ZDM. In other words, for sufficiently spin black holes and massive fields the $n=0$ mode simply decouples from the ZDM spectrum, being no longer described by Eq. (59). Consequently, the longest-living mode near extremality is no longer the $n=0$ mode (recall that we define the overtone numbers $n$ according to their behaviour in the Schwarzschild limit). We show this behavior in Fig. 8, where we draw contour plots of the continued fraction function $G(\omega)$ of Eq. (30). The plots are generated for a fixed $M\mu$ value near the critical mass $(M\mu)_{c}$, considering three distinct values of $a/M$: below, near, and above the critical spin $(a/M)_{c}$. In particular, we see that the $n=0$ mode, which is a ZDM for $a/M=0.99944<(a/M)_{c}$, becomes a DM when $a/M=0.99948>(a/M)_{c}$.

### v3

@@ II Scalar perturbations in Kerr Black Holes
$\partial_{u}\left[(1-u^{2})\partial_{u}S\right]+\left(\frac{m^{2}}{1-u^{2}}-c^{2}u^{2}-\Lambda\right)S=0,$ (6)
where $u=\cos\theta$, $c=a\sqrt{\omega^{2}-\mu^{2}}$, and $\Lambda=\lambda+2am\omega-a^{2}\omega^{2}$. It is also convenient to define
$\alpha=\sqrt{\omega^{2}-\mu^{2}},$ (7)
so that $c=a\alpha$.
@@
$R(r)\rightarrow\begin{cases}e^{-i(\omega-m\Omega_{+})r_{*}},\ \ \ &r\rightarrow r_{+}\ \ (r_{*}\rightarrow-\infty),\\ \frac{1}{r}e^{+i\alpha r_{*}},\ \ \ &r\rightarrow\infty\ \ (r_{*}\rightarrow\infty),\end{cases}$ (8)
@@ III.1 Boundary conditions and monodromy parameters
We note that $y_{t_{0},\pm}$ and $y_{\infty,\pm}$ form a basis of solutions near $z=t_{0}$ and $z=\infty$, respectively. Due to the Stokes’ phenomenon, however, the pair of solutions at the irregular singular point $z=\infty$ will only converge to an analytic function in the sector limited by the Stokes lines, which can be chosen to encompass the positive imaginary line in this case. In this region, the pair $y_{\infty,\pm}$ can be referred to as numerically satisfactory solutions after Ref. [68]. For the purposes of this article, we will assume that $\mathrm{Im}\,z>0$ in the following.
@@
$\mathsf{C}_{t}=\begin{pmatrix}e^{-\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{t_{0}}-e^{\tfrac{i\pi}{2}\eta}\zeta_{t_{0}}&-e^{-\tfrac{i\pi}{2}\eta}\zeta_{\infty}\zeta^{\prime}_{t_{0}}+e^{\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{\infty}\zeta_{t_{0}}\\ e^{-\tfrac{i\pi}{2}\eta}-e^{\tfrac{i\pi}{2}\eta}&-e^{-\tfrac{i\pi}{2}\eta}\zeta_{\infty}+e^{\tfrac{i\pi}{2}\eta}\zeta^{\prime}_{\infty}\end{pmatrix},$ (15)
@@
$\zeta_{\infty}=e^{-\frac{i\pi}{2}\sigma}\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma),$ (16a)
$\zeta^{\prime}_{\infty}=e^{\frac{i\pi}{2}\sigma}\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma),$ (16b)
@@
The QNM problem requires solutions of the radial equation to be purely ingoing at $z=t_{0}$ and purely outgoing at $z=\infty$. When $\mathrm{Re}\,\alpha>0$, this translates into $\mathsf{C}_{t}$ being a lower triangular matrix, and implies the following relation between the monodromy parameters
$e^{i\pi\eta}=\frac{\zeta_{\infty}\zeta^{\prime}_{t_{0}}}{\zeta^{\prime}_{\infty}\zeta_{t_{0}}}=e^{-i\pi\sigma}\frac{\sin\tfrac{\pi}{2}(\theta_{\star}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{\star}-\sigma)}\frac{\sin\tfrac{\pi}{2}(\theta_{t_{0}}+\theta_{0}+\sigma)}{\sin\tfrac{\pi}{2}(\theta_{t_{0}}+\theta_{0}-\sigma)}$
@@
We remark that $\rho$ and $\nu$ are canonically conjugate variables, in the sense that $\rho=(\pi i)^{-1}\partial_{\nu}{\cal W}$, just like the relationship between $\eta$ and $\sigma$ given in Eq. (23). Knowledge of $\eta$, together with $\sigma$, is required to obtain the transformation from $\{\sigma,\eta\}$ to $\{\nu,\rho\}$, allowing us to compute the accessory parameter $c_{t_{0}}$ from two different expansions, one in powers of $t_{0}$ and another in powers of $t_{0}^{-1}$. This provides not only for an independent numerical check of the eigenvalue calculation but also better behavior of the equations near extremality.
@@ III.2 Accessory parameter expansions via continued fractions
To compute the accessory parameter $c_{t_{0}}$ we employ a technique based on continued fractions. We consider Floquet type solutions of the CHE (10[73],
@@ III.3 The isomonodromic method for the QNM problem
$\theta_{0}=-m,\quad\theta_{t_{0}}=m,\quad\theta_{*}=0,\quad t_{0}=-4a\alpha$ (39a)
$t_{0}c_{t_{0}}=\lambda+2(1-m)a\alpha+a^{2}\alpha^{2}.$ (39b)
@@
For small $c=a\alpha$, we can derive an analytic expression for $\lambda=\lambda(\omega)$ by solving (41). Indeed, using the expansion (32) in powers of $t_{0}$, we find
$\begin{gathered}\lambda(c)=\ell(\ell+1)+\frac{1}{4}\bigg{(}\frac{2((\ell+1)^{2}-m^{2})(\ell+1)^{4}}{(2\ell+1)(\ell+1)^{3}(2\ell+3)}\\ \qquad\qquad-\frac{2(\ell^{2}-m^{2})\ell^{4}}{(2\ell-1)\ell^{3}(2\ell+1)}-1\bigg{)}c^{2}+{\mathcal{O}}(c^{3}).\end{gathered}$ (42)
For large $c$, a similar expression can be constructed from (37), with a suitable choice of $\nu$. In fact, the expression can be easily obtained from Eq. (4.7) in [41] by simply changing $a\omega$ to $c$. Since the expansion is not necessary for the remainder of this paper, we will leave out its explicit form.
@@
$z=2i\alpha(r-r_{-}).$ (43b)
Note that the region $r>r_{-}$ will map to the sector $\mathrm{Im}\,z>0$ provided $\mathrm{Re}\,\alpha>0$. As we will see below, this will become relevant as we increase $\mu$. The single monodromy parameters $\{\theta_{k}\}$ are expressed in terms of the physical parameters through
@@
$\theta_{\star}=\frac{2iM(2\omega^{2}-\mu^{2})}{\alpha}.$ (44c)
@@
$i\alpha(r_{-}-r_{+}-2iam+2i(a^{2}+r_{+}^{2})\omega)$
$+i\frac{M(2\omega^{2}-\mu^{2})}{\alpha}+\frac{M^{2}(2\omega^{2}-\mu^{2})^{2}}{\alpha^{2}}.$
@@
$g_{1}(\sigma,\eta,\omega)=0,$ (46)
provided $\mathrm{Re}\,\alpha>0$, which will be tested a posteriori. Additionally, substituting Eqs. (44) and (45) into the continued fraction equation (31), yields an implicit relation, denoted by $g_{2}$, between $\sigma$, $\omega$, and $\lambda$,
@@
Hence, in order determine the QNM frequencies through the isomonodromic method we need to determine the numerical solutions of the coupled system of algebraic equations (41), (46)-(48). In practice, we fix the black hole parameters $M$ and $a$ and the scalar field parameters $\mu$, $\ell$, $m$ and then determine the QNM frequency $\omega$ (and also the associated parameters $\lambda$, $\sigma$, and $\eta$). As explained before, there exists an infinite set of solutions $\omega_{n}$ indexed by the non-negative integer $n$. Our numerical routine uses the implementation of $\tau_{V}$ as a Fredholm determinant truncated at $N_{f}=64$ Fourier components. The continued fractions are truncated at $N_{c}=128$ and the root-finding algorithm we use is Muller’s method. The interested reader can find the details of our implementation of the isomonodromic method, based on the Julia language, in [38, 78]. Numerical codes and datasets are publicly available [79].
@@ IV.1 QNMs for $0\leq m<\ell$
Note that both the imaginary and the real parts of the QNM fundamental frequency converge to a finite value in the limit $a/M\rightarrow 1$. This finite value is the QNM frequency of the extremal black hole, computed using the RH map for the $\tau_{III}$-function defined in [38]. The $\tau_{III}$-function must be used instead of the $\tau_{V}$-function because the radial equation is doubly confluent when the black hole is extremal. We thus have found strong numerical evidence that the massive scalar QNM frequencies are continuous functions of the spin $a/M$ when $a/M=1$, just like the corresponding massless frequencies [48, 38].
@@ IV.2 QNMs for $\ell=m>0$
Regarding the results obtained in Figs. 1-3, we observe that the mass parameter impacts DMs and ZDMs differently in the extremal limit. The physical origin of ZDMs can be traced to the near-horizon geometry of nearly extreme black holes, whereas the origin of DMs is associated with the peaks of the potential barrier surrounding the black hole [46, 47]. Note that all modes in Fig. 1 are DMs and, therefore, are more strongly affected by changes in the mass in the extremal limit. In contrast, all modes in Fig. 3 are ZDMs and, as such, are less sensitive to variations in the mass parameter in the extremal limit. Finally, in Fig. 2, we observe both ZDMs and DMs. Notably, the DMs of Fig. 2 ($\mu M=0.4$ and $\mu M=0.5$) are as influenced by the mass of the scalar field as the DM modes of Fig. 1.
@@ V Zero-damping modes (ZDMs): Analytic expansion near extremality
in the near extremal regime. This expression mirrors the ones deduced in the literature for massless scalar fields around near extremal black holes [84, 85, 38] – see also Refs. [23, 86, 87, 88, 47, 89].
@@
We close this section with a discussion about the choice for the sign of $\sigma_{0}$ in Eq. (54a), which also determines the choice of the sign in the analytic expression (59) for the ZDMs near extremality. If $\sigma_{0}$ is real, which is expected for small $m$, we recall the analysis below Eq. (56) to justify that the signal of $\sigma_{0}$ must be chosen to enforce $\mathrm{Re}(\sigma_{0})>0$, which selects the positive root in (54). This seems to be the case only for the $\ell=m=1$ ZDM, as illustrated in Fig. 6, where we observe that the correction is of order $\mathcal{O}(\delta^{2},\delta\log\delta)$ to the real part of the frequencies $\omega_{k}$ for the first three ZDMs ($k=0$, $k=1$, and $k=2$). In particular, we see in the plots that the values of $\mathrm{Re}(M\omega_{k})$ for $k=1$ and $k=2$ intercept the corresponding value for the $k=0$ mode at $\delta\simeq 2.5\times 10^{-2}$ and at $\delta\simeq 1.5\times 10^{-2}$ respectively. For the imaginary part of the frequency $\mathrm{Im}(M\omega_{k})$, the linear $\delta$ term in (59) describes well the near extremal behaviour presented in Fig. 6. In Table 1, we compare the QNMs obtained through the numerical method detailed in Sec. III.3 with the frequencies calculated using the asymptotic expansion (59).
@@
Figure 7: Near extremal behavior of the first two ZDMs when $\ell=m=2$. The left and right panels show, respectively, the real and imaginary parts of the QNM frequencies for selected values of the scalar field mass.
@@ VI Continued fraction for large $t_{0}$ and contour plots
which is the continued fraction equation after the elimination of the dependence on the variables $\nu$ and $\lambda$. In order to gain insights on the QNM spectrum, we generate contour plots of the function $G(\omega)$ for complex values of the argument.
We start by checking whether the QNM frequencies calculated by finding the roots of $G$, which is based on large $t_{0}$ expansions, agrees with the results shown in Fig. 1 for $\ell=1$ and $m=0$, which were obtained with the isomonodromic method of Sec. III.3 that is based on power series of $t_{0}$. Table 2 presents the fundamental QNM frequency obtained by solving Eq. (60) for several masses $M\mu$ and spins $a/M$ when $\ell=1$ and $m=0$. We have confirmed that the values obtained from Fig. 1 are consistent with those listed in Table 2, with an agreement to order $10^{-10}$. Being DMs, as $a/M\rightarrow 1$, the real and imaginary parts of their frequencies converge to a nonzero finite value.
@@
We had previously identified the emergence of a bifurcation point for the $\ell=m=1$ fundamental QNM in the near extremal regime, as shown in Figs. 4 and 5. Beyond this bifurcation point, i.e. for $M\mu>(M\mu)_{c}\simeq 0.37049$, the $n=0$ QNM is a DM and not a ZDM. In other words, for sufficiently massive fields the $n=0$ mode simply decouples from the ZDM spectrum, being no longer described by Eq. (59). Consequently, the longest-living mode near extremality is no longer the $n=0$ mode (recall that we define the overtone numbers $n$ according to their behaviour in the Schwarzschild limit). We show this behavior in Fig. 8, where we draw contour plots of the continued fraction function $G(\omega)$ of Eq. (31). The plots are generated for a fixed $M\mu$ value near the critical mass $(M\mu)_{c}$, considering three distinct values of $a/M$: below, near, and above the critical spin $(a/M)_{c}$.